# The Role of Mixed-Language Documents for Multilingual Large Language Model Pretraining

### v1

@@ 1 Introduction
Large language models (LLMs), when pretrained on web-collected data from different language sources, exhibit remarkable emergent capabilities in cross-lingual understanding despite not being pretrained using multilingual-specific objectives (mBERT; Achiam2023GPT4TR; Yang2024Qwen25TR). Existing research attributes this behaviour not only to a sufficient amount of data from different languages, but also to specific documents where multiple languages co-occur in the same context (Chaudhary2020DICTMLMIM; Chi2020InfoXLMAI; Wang2025InvestigatingAS). Motivated by this observation, multilingual pretraining strategies have often incorporated multilingual data, under the hypothesis that mixed-language exposure uniformly benefits cross-lingual tasks (Yoo2024CodeSwitchingCL; Wang2025InvestigatingAS).
@@
However, the high cost of pretraining and large-scale pretraining data classfication has constrained the scope of existing explorations of the role of multilingual data. Studies that rely on continual pretraining typically build on models that may have already been exposed to related data during pretraining, which makes the role of multilingual data more difficult to disentangle. Among the few works that investigate multilingual data at the pretraining stage, existing studies cherry; Qorib2025JustGP; Wang2025InvestigatingAS do not provide a systematic analysis of its role, but instead focus on specific settings or mechanisms. To this end, we aim to conduct a thorough analysis of the pretraining corpus and design a controlled pretraining setup to explicitly reveal the role of multilingual data.
@@ 2 Related Work
Multilingual data is widely assumed to drive cross-lingual capabilities in multilingual models. Parallel corpora (sentence-aligned translations) are well-known to be essential for machine translation (Brown1993TheMO), enabling multilingual MT systems to bridge high- and low-resource languages (Johnson2016GooglesMN; Fan2020BeyondEM; team2022NoLL). Beyond parallel data, naturally occurring code-switching, where languages alternate within the same discourse, has also attracted attention as a potential mechanism for cross-lingual alignment. Prior work shows that using code-switching for data augmentation can improve zero-shot transfer during finetuning (Qin2020CoSDAMLMC), and that curriculum learning with code-switching enhances transfer to low-resource languages (Yoo2024CodeSwitchingCL). These results have motivated practitioners to incorporate multilingual content under the assumption that mixed-language exposure benefits cross-lingual tasks (Qorib2025JustGP).
However, most existing studies focus on finetuning or continued training, which remains limited because models may have already been exposed to similar data during the pretraining stage. Among the few works that investigate multilingual data in the pretraining stage, most focus on specific aspects rather than systematically studying its role: one primarily characterizes incidental bilingualism in existing corpora (cherry), another uses generated data to study curriculum learning effects (Qorib2025JustGP), and a third explores synthetic code-switching for cross-lingual transfer (Wang2025InvestigatingAS). This leaves a gap in the understanding of the role of multilingual data during pretraining, particularly regarding the differential contributions of parallel versus code-switching data. We aim to fill this gap by systematically ablating different bilingual data types and studying their impact on multilingual LLMs.
@@ 3 MonoWeb Pretraining Data
To study the heterogeneous role of bilingual data, we first construct a multilingual corpus by sampling 60B tokens per language from FineWeb-Edu (lozhkov2024fineweb-edu) (English) and FineWeb2 (fineweb2) (German, Spanish, French), totaling 240B tokens. We then perform a systematic characterization of bilingual documents in this corpus, focusing on English-paired bilingual content (en-de, en-es, en-fr) as English serves as the current dominant lingua franca for cross-lingual scenarios.
@@ 3.1 Bilingual Data Identification
We first detect candidate bilingual documents using language-level entropy as a proxy for language mixing. For each document, we perform sentence segmentation using NLTK (nltk) and apply fastText language identification (fasttext) to compute language confidence scores for each sentence. Taking English-French as an example, for each sentence $s_{i}$ with length $l_{i}$, fastText outputs confidence scores for English ($p_{i}^{\text{en}}$) and French ($p_{i}^{\text{fr}}$). We then compute a document-level language distribution by aggregating sentence-level scores weighted by sentence length:
@@
To distinguish different types of bilingual relationships, we employ Llama-3.3-70B-Instruct (Llama3) for a two-step classification process, whose reliability has been validated through human evaluation. First, the model verifies whether each candidate is genuinely bilingual, which aims to filter out the false negatives included by entropy-based filtering. We consider the resulting set of documents after this step as the final verified bilingual documents, which consists of approximately 2% of the entire corpus. Second, based on the semantic relationship of contents in different languages, the verified bilingual documents are classified into one of the three categories:
@@ 4.2 Model Architecture and Training
We train decoder-only transformer models with 1.35B parameters using the Llama-2 tokenizer (Touvron2023Llama2O) (32K vocabulary). The architecture consists of 24 layers with a 2048 hidden dimensions, 16 attention heads, and a 2048 context length. All models are trained for 34K steps (1̃43B tokens) with a batch size of 2,048 using the AdamW optimizer (Loshchilov2017DecoupledWD) and a 6e-4 learning rate, including 2,000 warmup steps followed by constant decay. We set weight decay to 0.1, apply gradient clipping at 1.0, and use Adam betas of 0.9 and 0.95. Training is performed with Megatron-LM (megatron-lm) on 128 NVIDIA GH200 GPUs and takes about 8 hours per model.
@@ 4.3 Downstream Evaluation Suite
All tasks are evaluated using the lm-evaluation-harness (eval-harness), along with five-shot prompting and default configurations.
@@
We evaluate translation quality on standard benchmarks for all three language pairs: wmt16 en-de (WMT16), wmt14 en-fr (wmt14), and flores-101 en-es (FLORES101), testing both translation directions and report BLEU scores (bleu) separately for each direction.
@@
We evaluate cross-lingual question answering using two complementary benchmarks. (1) For XQuAD (XQuAD), we adapt the dataset by placing the context in language L1 and both the question and answer in language L2, allowing us to assess the model’s ability to generate answers across languages. (2) For MLQA (lewis2019mlqa), we follow the original setup, where the context and answer are in language L1 and the question is in language L2, which primarily evaluates the model’s ability to retrieve information across languages. We report Exact Match scores for all language pairs.
@@
We evaluate models on a suite of benchmarks covering both cross-lingual understanding and reasoning abilities. For cross-lingual natural language understanding, we use XNLI (XNLI) and PAWS-X (paws-x) to assess whether bilingual data improves the transfer of inference and paraphrase recognition skills. For reasoning tasks, we include HellaSwag (Hellswag; Hellswag_arc_multilingual) for commonsense reasoning, ARC (Hellswag_arc_multilingual; arc) for knowledge-intensive reasoning, TruthfulQA (truthfulqa; Hellswag_arc_multilingual) for factual consistency, and additionally XStoryCloze (xstorycloze) (en, es) and XWinograd (xwinograd) (en, fr) for narrative comprehension and coreference resolution. We report accuracy for all tasks.
@@ 5.4 Explaining the Asymmetry: Why MT Collapses but Reasoning Persists
Removing bilingual data causes a severe collapse in machine translation, while cross-lingual reasoning and understanding tasks remain largely unaffected. To explain this phenomenon, we analyze how bilingual data removal impacts cross-lingual alignment at different linguistic granularities. Specifically, we measure alignment across layers using Precision@1 (P@1) computed with cosine similarity for sentence representations (3,000 WMT parallel sentences) and word representations (2,000 MUSE MUSE pairs). As shown in Table 8, we observe a stark divergence: while MonoWeb preserves robust sentence-level alignment ($2\%$ drop from FineWeb), it suffers a sharp 13–21% degradation in lexical-level alignment. This suggests that monolingual pretraining is sufficient to align sentence-level semantics, supporting cross-lingual understanding and reasoning, but fails to establish the fine-grained lexical correspondences required for accurate translation.
@@ 7 Limitations
8 Acknowledgements
We thank Jiayi Wang for her helpful and constructive comments.

### v2

@@ 1 Introduction
Large language models (LLMs), when pretrained on web-collected data from different language sources, exhibit remarkable emergent capabilities in cross-lingual understanding despite not being pretrained using multilingual-specific objectives (Devlin et al., 2019; Achiam et al., 2023; Yang et al., 2024). Existing research attributes this behaviour not only to a sufficient amount of data from different languages, but also to specific documents where multiple languages co-occur in the same context (Chaudhary et al., 2020; Chi et al., 2020; Wang et al., 2025). Motivated by this observation, multilingual pretraining strategies have often incorporated multilingual data, under the hypothesis that mixed-language exposure uniformly benefits cross-lingual tasks (Yoo et al., 2024; Wang et al., 2025).
@@
However, the high cost of pretraining and large-scale pretraining data classfication has constrained the scope of existing explorations of the role of multilingual data. Studies that rely on continual pretraining typically build on models that may have already been exposed to related data during pretraining, which makes the role of multilingual data more difficult to disentangle. Among the few works that investigate multilingual data at the pretraining stage, existing studies Briakou et al. (2023); Qorib et al. (2025); Wang et al. (2025) do not provide a systematic analysis of its role, but instead focus on specific settings or mechanisms. To this end, we aim to conduct a thorough analysis of the pretraining corpus and design a controlled pretraining setup to explicitly reveal the role of multilingual data.
@@ 2 Related Work
Multilingual data is widely assumed to drive cross-lingual capabilities in multilingual models. Parallel corpora (sentence-aligned translations) are well-known to be essential for machine translation (Brown et al., 1993), enabling multilingual MT systems to bridge high- and low-resource languages (Johnson et al., 2016; Fan et al., 2020; team et al., 2022). Beyond parallel data, naturally occurring code-switching, where languages alternate within the same discourse, has also attracted attention as a potential mechanism for cross-lingual alignment. Prior work shows that using code-switching for data augmentation can improve zero-shot transfer during finetuning (Qin et al., 2020), and that curriculum learning with code-switching enhances transfer to low-resource languages (Yoo et al., 2024). These results have motivated practitioners to incorporate multilingual content under the assumption that mixed-language exposure benefits cross-lingual tasks (Qorib et al., 2025).
However, most existing studies focus on finetuning or continued training, which remains limited because models may have already been exposed to similar data during the pretraining stage. Among the few works that investigate multilingual data in the pretraining stage, most focus on specific aspects rather than systematically studying its role: one primarily characterizes incidental bilingualism in existing corpora (Briakou et al., 2023), another uses generated data to study curriculum learning effects (Qorib et al., 2025), and a third explores synthetic code-switching for cross-lingual transfer (Wang et al., 2025). This leaves a gap in the understanding of the role of multilingual data during pretraining, particularly regarding the differential contributions of parallel versus code-switching data. We aim to fill this gap by systematically ablating different bilingual data types and studying their impact on multilingual LLMs.
@@ 3 MonoWeb Pretraining Data
To study the heterogeneous role of bilingual data, we first construct a multilingual corpus by sampling 60B tokens per language from FineWeb-Edu (Lozhkov et al., 2024) (English) and FineWeb2 (Penedo et al., 2025) (German, Spanish, French), totaling 240B tokens. We then perform a systematic characterization of bilingual documents in this corpus, focusing on English-paired bilingual content (en-de, en-es, en-fr) as English serves as the current dominant lingua franca for cross-lingual scenarios.
@@ 3.1 Bilingual Data Identification
We first detect candidate bilingual documents using language-level entropy as a proxy for language mixing. For each document, we perform sentence segmentation using NLTK (Bird et al., 2009) and apply fastText language identification (Joulin et al., 2016) to compute language confidence scores for each sentence. Taking English-French as an example, for each sentence $s_{i}$ with length $l_{i}$, fastText outputs confidence scores for English ($p_{i}^{\text{en}}$) and French ($p_{i}^{\text{fr}}$). We then compute a document-level language distribution by aggregating sentence-level scores weighted by sentence length:
@@
To distinguish different types of bilingual relationships, we employ Llama-3.3-70B-Instruct (Dubey et al., 2024) for a two-step classification process, whose reliability has been validated through human evaluation. First, the model verifies whether each candidate is genuinely bilingual, which aims to filter out the false negatives included by entropy-based filtering. We consider the resulting set of documents after this step as the final verified bilingual documents, which consists of approximately 2% of the entire corpus. Second, based on the semantic relationship of contents in different languages, the verified bilingual documents are classified into one of the three categories:
@@ 4.2 Model Architecture and Training
We train decoder-only transformer models with 1.35B parameters using the Llama-2 tokenizer (Touvron et al., 2023) (32K vocabulary). The architecture consists of 24 layers with a 2048 hidden dimensions, 16 attention heads, and a 2048 context length. All models are trained for 34K steps (1̃43B tokens) with a batch size of 2,048 using the AdamW optimizer (Loshchilov and Hutter, 2017) and a 6e-4 learning rate, including 2,000 warmup steps followed by constant decay. We set weight decay to 0.1, apply gradient clipping at 1.0, and use Adam betas of 0.9 and 0.95. Training is performed with Megatron-LM (Shoeybi et al., 2019) and takes about 6,144 A100 GPU hours per model.
@@ 4.3 Downstream Evaluation Suite
All tasks are evaluated using the lm-evaluation-harness (Gao et al., 2024), along with five-shot prompting and default configurations.
@@
We evaluate translation quality on standard benchmarks for all three language pairs: wmt16 en-de (Bojar et al., 2016), wmt14 en-fr (Bojar et al., 2014), and flores-101 en-es (Goyal et al., 2021), testing both translation directions and report BLEU scores (Papineni et al., 2002) separately for each direction.
@@
We evaluate cross-lingual question answering using two complementary benchmarks. (1) For XQuAD (Artetxe et al., 2019), we adapt the dataset by placing the context in language L1 and both the question and answer in language L2, allowing us to assess the model’s ability to generate answers across languages. (2) For MLQA (Lewis et al., 2019), we follow the original setup, where the context and answer are in language L1 and the question is in language L2, which primarily evaluates the model’s ability to retrieve information across languages. We report Exact Match scores for all language pairs.
@@
We evaluate models on a suite of benchmarks covering both cross-lingual understanding and reasoning abilities. For cross-lingual natural language understanding, we use XNLI (Conneau et al., 2018) and PAWS-X (Yang et al., 2019) to assess whether bilingual data improves the transfer of inference and paraphrase recognition skills. For reasoning tasks, we include HellaSwag (Zellers et al., 2019; Dac Lai et al., 2023) for commonsense reasoning, ARC (Dac Lai et al., 2023; Clark et al., 2018) for knowledge-intensive reasoning, TruthfulQA (Lin et al., 2022; Dac Lai et al., 2023) for factual consistency, and additionally XStoryCloze (Lin et al., 2021) (en, es) and XWinograd (Tikhonov and Ryabinin, 2021) (en, fr) for narrative comprehension and coreference resolution. We report accuracy for all tasks.
@@ 5.4 Explaining the Asymmetry: Why MT Collapses but Reasoning Persists
Removing bilingual data causes a severe collapse in machine translation, while cross-lingual reasoning and understanding tasks remain largely unaffected. To explain this phenomenon, we analyze how bilingual data removal impacts cross-lingual alignment at different linguistic granularities. Specifically, we measure alignment across layers using Precision@1 (P@1) computed with cosine similarity for sentence representations (3,000 WMT parallel sentences) and word representations (2,000 MUSE Conneau et al. (2017) pairs). As shown in Table 8, we observe a stark divergence: while MonoWeb preserves robust sentence-level alignment ($2\%$ drop from FineWeb), it suffers a sharp 13–21% degradation in lexical-level alignment. This suggests that monolingual pretraining is sufficient to align sentence-level semantics, supporting cross-lingual understanding and reasoning, but fails to establish the fine-grained lexical correspondences required for accurate translation.